OpenStudy (lilsis76):

graph each point in a polar coordinate system, then convert the given polar coordinate to rectangular coordinates. (-1, pi/8) this is all my homework is and i have other ones where i have to reverse the coordinates from recangle to polar

OpenStudy (anonymous):

Polar coords give you $$(r,\theta)$$ so that on the axis you draw something like  you can then convert that point to an (x,y) rectangular coord. by using trig and making a triangle. $$y=\sin(\theta) \\ x = \cos(\theta) $$

OpenStudy (anonymous):



OpenStudy (lilsis76):

then where does the pi/8 land? i cant find that

OpenStudy (anonymous):

radian notation can be converted as follows: 0pi = 0 degrees pi = 180 degrees 2pi = 360 degrees so pi/8 = 180/8 = 22.5 degrees. first quadrant

OpenStudy (lilsis76):

what the heck?! haha

OpenStudy (lilsis76):

i didnt think about the 180 over 8

OpenStudy (lilsis76):

dang it. now im lost :/ i dont know what to do next

OpenStudy (anonymous):

so to get the (x,y) you use (cos(angle), sin(angle)) so for this one \[\bigg(\cos\left(\frac{\pi}{8}\right),\sin\left(\frac{\pi}{8}\right)\bigg)\]